# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v1

@@ Abstract
While neural networks have demonstrated impressive performance across various tasks, accurately quantifying uncertainty in their predictions is essential to ensure their trustworthiness and enable widespread adoption in critical systems. Several Bayesian uncertainty quantification (UQ) methods exist that are either cheap or reliable, but not both. We propose a post-hoc, sampling-based uncertainty quantification (UQ) method for overparameterized networks at the end of training. Our approach constructs efficient and meaningful deep ensembles by employing a (stochastic) gradient-descent sampling process on appropriately linearized networks. We demonstrate that our method effectively approximates the posterior of a Gaussian Process using the empirical Neural Tangent Kernel. Through a series of numerical experiments, we show that our method not only outperforms competing approaches in computational efficiency–often reducing costs by multiple factors–but also maintains state-of-the-art performance across a variety of UQ metrics for both regression and classification tasks.
Bayesian Deep Learning, Neural Tangent Kernel, Uncertainty Quantification, ICML
[…]
@@ 1 Introduction
Figure 1: Comparison of various Bayesian UQ methods (see Section 2) on a 1-layer MLP, trained on the data (red) lying on $y=x^{3}$ (black), with Gaussian noise added. The methods’ mean predictors (blue) ±$\pm 3\sigma$ (green) are shown, where $\sigma^{2}$ is the variance estimated via each method. Clearly, VI, SWAG and LA all perform poorly on this task, while DE performs better, and LLA and NUQLS perform well.
@@
An array of uncertainty quantification methods for NNs exist, each with benefits and drawbacks. Frequentist statistical methods, such as conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002; Lei & Wasserman, 2014), create prediction intervals through parameter estimation and probability distributions on the data. While currently considered state-of-the-art, the main drawback is that conformal prediction is data-hungry, requiring a large hold-out set. In contrast, parametric Bayesian methods model the uncertainty in the parameters, providing conditional predictive probabilities for each function output.
@@
Several works (Altieri et al., 2024; Dadalto et al., 2023; Granese et al., 2021) employ the feature-space representations of the network to quantify the uncertainty of test predictions, through kernel densities (Kotelevskii et al., 2022), Gaussian Discriminant analysis (Mukhoti et al., 2023), non-constant mapping functions (Tagasovska & Lopez-Paz, 2019), etc. The interpretation of uncertainty in these works is related to risk of misclassification; these methods generally perform very well at detecting OoD points.While both conformal predictions and the feature-space works just mentioned are important, we limit our scope to Bayesian methods in this paper. Though Bayesian methods can suffer from prior misspecification (Masegosa, 2020) and computational burdens, the predictive and posterior distributions of a model are very natural frameworks for quantifying the uncertainty and spread of possible values of the model.
@@
    We present a Monte-Carlo sampling based UQ method to approximate the predictive distribution of NNs, called Neural Uncertainty Quantification by Linearized Sampling (NUQLS). Our method is lightweight, post-hoc, numerically stable, and embarrassingly parallel.
@@
    Under certain assumptions, we show that NUQLS converges to the predictive distribution of a GP with an empirical NTK kernel, providing a novel perspective on the connection between NNs, GPs, and the NTK.
@@
    On various ML problems, we show that NUQLS performs as well as or better than leading UQ methods, is less computationally expensive than deep ensemsble, and scales to large image classification tasks.
@@
Throughout the paper, we denote scalars, vectors, and matrices as lower-case, bold lower-case, and bold upper-case letters, e.g., $c$, $\bm{\theta}$ and ${{\mathbf{K}}}$, respectively. For two vectors ${\bf v}\in^{p}$ and ${\bf w}\in^{p}$, their Euclidean inner product is denoted as $\left\langle{\bf v},{\bf w}\right\rangle={\bf v}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{\bf w}$. We primarily consider supervised learning, which involves a function ${\bf f}:^{d}\times^{p}\to^{c}$, assumed sufficiently smooth in its parameters $\bm{\theta}\in^{p}$, a training dataset $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}=\{{\bf x}_{i},{\bf y}_{i}\}_{i=1}^{n}\subset^{d}\times^{c}$, and a loss function $\ell:^{c}\times^{c}\to[0,\infty)$.
@@ 2 Background
[…]
Proposed in (Gal & Ghahramani, 2016), Monte Carlo Dropout (MC-Dropout) takes a trained NN, and uses the dropout regularization technique at test time to sample $S$ sub-networks, $\left\{{\bf f}(\bm{\theta}_{s},{\bf x})\right\}_{s=1}^{S}$ as an MC approximation of the predictive distribution. MC-Dropout is an inexpensive method, yet it is unlikely to converge to the true posterior, and is erroneously multi-modal (Folgoc et al., 2021).
@@
A generalized Gauss-Newton (GGN) approximation to the Hessian ensures a positive-definite Fisher approximation (Khan et al., 2019). Moreover, to scale, the LA requires reduction to a subset of parameters or further approximations of the covariance structure, such as only forming the posterior over parameters in the last dense layer of a NN (last-layer), or using Kronecker-Factored Approximate Curvature (KFAC) (Martens & Grosse, 2015).
We can evaluate the posterior and predictive distribution in the LA using the linearization of ${\bf f}({\bf x},\bm{\theta})$ around the MAP solution. This approach is known as the Linearized Laplace Approximation (LLA) and typically delivers better performance than LA (Immer et al., 2021). Recent work (Antorán et al., 2022; Ortega et al., 2023) has enabled LLA to become more scalable for larger models and datasets.
@@
for ${\bf y}\triangleq[y_{1},\dots,y_{n}]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}$ and $\bm{\mu}(\mathcal{X})\triangleq[\mu({\bf x}_{1}),\dots,\mu({\bf x}_{n})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}$.
GPs can yield impressive predictive results when a suitable kernel is chosen (Rasmussen, 1997). However, forming the kernel and solving linear systems makes GP computations intractable for large datasets. Approximations such as sparse variational inference (Titsias, 2009), Nyström methods (Martinsson & Tropp, 2020), and other subspace approximations (Gardner et al., 2018) can alleviate the computational burden; however, these approximations often result in a significant decline in predictive performance.
@@ 3 NUQLS
We now present Neural Uncertainty Quantification by Linearized Sampling (NUQLS), our post-hoc sampling method for quantifying the uncertainty of a trained NN. We begin by presenting the motivation and a high-level overview of our method. Subsequently, we provide theoretical justification, demonstrating that, under specific conditions, the NUQLS samples represent draws from the approximate posterior of the neural network, which is equivalent to a Gaussian process defined by the NTK.
@@ 3.1 Motivation and High-level Overview
${\bf f}(\bm{\theta},{\bf x})$ $\approx\widetilde{{\bf f}}(\bm{\theta},{\bf x})\triangleq{\bf f}(\widehat{\bm{\theta}},{\bf x})+{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x})(\bm{\theta}-\widehat{\bm{\theta}}),$ (2)
where ${\mathbf{J}}(.,{\bf x})=[\partial{\bf f}(.,{\bf x})/\partial\bm{\theta}]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\in^{c\times p}$ is the Jacobian of ${\bf f}$. Using the linear approximation 2, we consider
@@
In overparameterized settings, 3 may have infinitely many solutions. To identify these solutions and create our ensemble of linear predictors, we employ (stochastic) gradient descent, initialized at zero-mean isotropic Gaussian perturbations of the trained parameter, $\widehat{\bm{\theta}}$. The pseudo-code for this algorithm is provided in Algorithm 1.
Note that while the training cost for a linearised network is only slightly higher per epoch compared to standard NN training, each network in the NUQLS ensemble is initialized in a neighborhood of a local minimum of the original NN. As a result, NUQLS often requires significantly fewer epochs to converge, leading to an order-of-magnitude computational speedup relative to DE (see Tables 1, 2 and 6 for wall-clock time comparisons.).
Algorithm 1 NUQLS
  Input: Number of realizations, $S$, NN weights, $\widehat{\bm{\theta}}$.
  for $s=1$ : $S$ do
     $\bm{\theta}^{\star}_{s}\leftarrow$ Run (stochastic) gradient descent from $\bm{\theta}_{0,s}$ to (approximately) solve 3 and obtain $\bm{\theta}^{\star}_{s}$
  end for
  return  $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},.)\}_{s=1}^{S}$
For a given test point ${{{\bf x}}^{\star}}$, the mean prediction and the associated uncertainty can be computed using $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x}^{\star})\}_{s=1}^{S}$.
@@ 3.2 Theoretical Analysis
We now establish the key property of Algorithm 1: under mild conditions, NUQLS generates samples from the approximate posterior of the neural network, which in many cases corresponds to a Gaussian process defined by the NTK. Proofs are provided in Appendix A.
Suppose $\bm{\theta}^{\ddagger}$ is any solution to 3. Using $\bm{\theta}^{\ddagger}$, one can construct a family of solutions to 3 as
@@
where $\widehat{\bm{\theta}}$ is the parameters of the trained NN and ${\mathbf{J}}_{\mathcal{X}}=[{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}(\widehat{\bm{\theta}},{\bf x}_{1})\;\ldots\;{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}(\widehat{\bm{\theta}},{\bf x}_{n})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\in^{nc\times p}$. Note that the second term in 4 consists of all vectors in the null space of ${\mathbf{J}}_{\mathcal{X}}$. Since any such $\bm{\theta}^{\ddagger}$ can be decomposed as the direct sum of components in the null space of ${\mathbf{J}}_{\mathcal{X}}$ and its orthogonal complement, the family of solutions in 4 depends on the choice of $\bm{\theta}^{\ddagger}$. However, under certain assumptions, we can ensure that the representation 4 is uniquely determined, i.e., $\bm{\theta}^{\ddagger}$ can be taken as the unique solution to 3 that is orthogonal to the null space of ${\mathbf{J}}_{\mathcal{X}}$. More precisely, we can show that, under these assumptions on the loss, $\bm{\theta}^{\ddagger}$ in 4 can be taken as the unique solution to
@@
  • strictly convex in its first argument, and a solution to 3 exists.
The problem 5 admits a unique solution.
As it turns out, any solution of the form 4 can be efficiently obtained using (stochastic) gradient descent.
@@
Consider the optimization problem 3 and assume ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank.
  • (Gradient Descent) Suppose $\ell({\bf f},{\bf y})$ is strictly convex with locally Lipschitz continuous gradient, both with respect to ${\bf f}$, and the problem 3 admits a solution. Gradient descent, initialized at ${\bf z}$ and with appropriate learning rate, converges to 4.
  • (Stochastic Gradient Descent) Suppose $\ell({\bf f},{\bf y})$ is strongly convex with Lipschitz continuous gradient, both with respect to ${\bf f}$, and any solution to the problem 3 is interpolating. Stochastic gradient descent, initialized at ${\bf z}$ and with small enough learning rate, converges to 4 with probability one.
We note that the local smoothness requirement in the first part of Theorem 3.2 is a relatively mild assumption; for example, it holds if we simply assume that $\ell$ is twice continuously differentiable. Also, the full-row rank assumption on ${\mathbf{J}}_{\mathcal{X}}$ in the second part of Theorem 3.2 is reasonable for highly over-parameterized networks; e.g., see (Liu et al., 2022).
Now, suppose ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank and the assumption of Lemma 3.1 holds, ensuring the existence of the unique solution $\bm{\theta}^{\ddagger}$ to 5. Noting ${\textnormal{Range}}\left({\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)={\textnormal{Range}}\left({\mathbf{J}}_{\mathcal{X}}^{\dagger}\right)$, we can write 4 as
@@
where ${\bf w}$ is some vector for which $\bm{\theta}^{\ddagger}={\mathbf{J}}_{\mathcal{X}}^{\dagger}{\bf w}$, and ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}\triangleq{\mathbf{J}}_{\mathcal{X}}{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}_{\mathcal{X}}={{\mathbf{K}}}_{\widehat{\bm{\theta}}}(\mathcal{X},\mathcal{X})\in^{nc\times nc}$ is the Gram matrix of the empirical NTK 1 on the training data $\mathcal{X}$. Setting ${\bf z}=\widehat{\bm{\theta}}-{\bf z}_{0}$ for some ${\bf z}_{0}$, we get

### v2

@@ Abstract
While neural networks have demonstrated impressive performance across various tasks, accurately quantifying uncertainty in their predictions is essential to ensure their trustworthiness and enable widespread adoption in critical systems. Several Bayesian uncertainty quantification (UQ) methods exist that are either cheap or reliable, but not both. We propose a post-hoc, sampling-based UQ method for overparameterized networks at the end of training. Our approach constructs efficient and meaningful deep ensembles by employing a (stochastic) gradient-descent sampling process on appropriately linearized networks. We demonstrate that our method effectively approximates the posterior of a Gaussian Process using the empirical Neural Tangent Kernel. Through a series of numerical experiments, we show that our method not only outperforms competing approaches in computational efficiency–often reducing costs by multiple factors–but also maintains state-of-the-art performance across a variety of UQ metrics for both regression and classification tasks.
@@ 1 Introduction
An array of uncertainty quantification methods for NNs exist, each with benefits and drawbacks. Frequentist statistical methods, such as conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002; Lei & Wasserman, 2014), create prediction intervals through parameter estimation and probability distributions on the data. While currently considered state-of-the-art, the main drawback is that conformal prediction is data-hungry, requiring a large hold-out set. Several works Altieri et al. (2024); Dadalto et al. (2023); Granese et al. (2021) employ the feature-space representations of the network to quantify the uncertainty of test predictions, through kernel densities (Kotelevskii et al., 2022), Gaussian Discriminant analysis (Mukhoti et al., 2023), non-constant mapping functions (Tagasovska & Lopez-Paz, 2019), etc. The interpretation of uncertainty in these works is related to risk of misclassification; these methods generally perform very well at detecting OoD points. While both conformal predictions and feature-space works are important, we limit our scope to Bayesian methods in this paper. Though Bayesian methods can suffer from prior misspecification (Masegosa, 2020) and computational burdens, the predictive and posterior distributions of a model are very natural frameworks for quantifying the uncertainty and spread of possible values of the model.
Figure 1: Comparison of various Bayesian UQ methods (see Section˜2) on a 1-layer MLP, trained on the data (red) lying on $y=x^{3}$ (black), with Gaussian noise added. The methods’ mean predictors (blue) $\pm 3\sigma$ (green) are shown, where $\sigma^{2}$ is the variance estimated via each method. We see that NUQLS performs well on this task.
@@
    In Section˜3.1, we present a Monte-Carlo sampling based UQ method to approximate the predictive distribution of NNs, called Neural Uncertainty Quantification by Linearized Sampling (NUQLS). Our method is lightweight, post-hoc, numerically stable, and embarrassingly parallel.
@@
    Under certain assumptions, Section˜3.2 details the convergence of NUQLS to the predictive distribution of a GP with an empirical NTK kernel, providing a novel perspective on the connection between NNs, GPs, and the NTK.
@@
    On various ML problems in Section˜4, we show that NUQLS performs as well as or better than leading UQ methods, is less computationally expensive than deep ensemble, and scales to large image classification tasks.
  4. 4.
    In Section˜4.4 we introduce a novel metric for evaluating the quality of UQ methods for classification tasks. This metric more directly measures the quality of UQ methods than existing UQ metrics.
Remark 1.1 (Necessity of Contribution 4.).
Evaluating the quality of UQ methods is non-trivial. There exists in the literature a lack of a suitable framework for evaluating the quality of UQ methods in classification settings. Common metrics such as Negative Log-Likelihood (NLL), Expected Calibration Error (ECE) (Naeini et al., 2015) and Area Under the Curve of the Receiver Operating Characteristic (AUCROC) are actually metrics for prediction quality, or are based on flawed surrogates for uncertainty. Further, it was shown in Abe et al. (2022) that the goal of Bayesian methods should not be to improve predictive ability; performance gains will be attained more economically by choosing a larger model class. Instead, Bayesian methods should seek to accurately quantify the uncertainty of a model, by computing the predictive variance. As uncertainty is not a measurement that can easily be shown to be well-calibrated, one requires a more qualitative approach to evaluating the performance of a UQ model. This is the motivation for the graphical technique we introduce in Section˜4.4 for comparing the quality of UQ estimates for multi-class classification. For an in-depth discussion of these points, please see Appendix˜C.
@@
[…]
@@ 2 Background
Parametric Bayesian methods admit access to a distribution over predictions ${\bf f}(\bm{\theta},{\bf x}^{\star})$ for unseen test points $\bf x$$^{\star}$, through the posterior distribution $p(\bm{\theta}|\mathcal{D})\propto p(\bm{\theta})p(\mathcal{D}|\bm{\theta})$, and the predictive distribution $p({\bf y}^{\star}|{\bf x}^{\star},\mathcal{D})=\int p({\bf y}^{\star}|{\bf f}(\bm{\theta},{\bf x}^{\star}))p(\bm{\theta}|\mathcal{D})d\bm{\theta}$, where $p(\bm{\theta})$ is the prior, and $p(\mathcal{D}|\bm{\theta})$ is the likelihood function evaluated on the training data. Both the posterior and the predictive distributions are computationally intractable in all but the simplest cases. In our setting, their calculation involves very high-dimensional integrals, which we can approximate through a Monte Carlo (MC) approximation $p({\bf y}^{\star}|{\bf x}^{\star},\mathcal{D})\approx 1/S\sum_{s}p({\bf y}^{\star}|{\bf f}(\bm{\theta}_{s},{\bf x}^{\star}))$ for $\bm{\theta}_{s}\sim q(\bm{\theta})$, where $q(\bm{\theta})$ is an approximation to the posterior distribution. The effectiveness of classical Markov Chain Monte Carlo (MCMC) methods in posterior sampling diminishes in this setting due to the curse of dimensionality, limiting the tractable techniques available with theoretical guarantees. This limitation necessitates coarser approximations for estimating the posterior $q(\bm{\theta})$, leading to the emergence of the following Bayesian methods for posterior approximation.
Proposed in (Gal & Ghahramani, 2016), Monte Carlo Dropout (MC-Dropout) takes a trained NN, and uses the dropout regularization technique at test time to sample $S$ sub-networks, $\left\{{\bf f}(\bm{\theta}_{s},{\bf x})\right\}_{s=1:S}$ as an MC approximation of the predictive distribution. MC-Dropout is an inexpensive method, yet it is unlikely to converge to the true posterior, and is erroneously multi-modal (Folgoc et al., 2021).
@@
We can evaluate the posterior and predictive distribution in the LA using the linearization of ${\bf f}({\bf x},\bm{\theta})$ around the MAP solution. This approach is known as the Linearized Laplace Approximation (LLA) and typically delivers better performance than LA (Immer et al., 2021). LLA generally requires reduction to a subset of parameters or approximations of the covariance structure (Martens & Grosse, 2015) to scale, at the cost of performance. Recent work (Antorán et al., 2022; Ortega et al., 2023) has enabled LLA to become more scalable with better performance for larger models and datasets.
@@
for ${\bf y}\triangleq[y_{1},\dots,y_{n}]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}$ and $\bm{\mu}(\mathcal{X})\triangleq[\mu({\bf x}_{1}),\dots,\mu({\bf x}_{n})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}$. GPs can yield impressive predictive results when a suitable kernel is chosen (Rasmussen, 1997). However, forming the kernel and solving linear systems makes GP computations intractable for large datasets. Approximations such as sparse variational inference (Titsias, 2009), Nyström methods (Martinsson & Tropp, 2020), and other subspace approximations (Gardner et al., 2018) can alleviate the computational burden; however, these approximations often result in a significant decline in predictive performance.
@@
Related Works
Our method NUQLS shares notable similarities with, and exhibits distinct differences from, several prior works, such as Sampling-LLA, Bayesian Deep Ensembles and local ensembles. Due to the breadth and depth of this discussion, we relegate the related works discussion to Appendix˜B.
@@ 3 NUQLS
Algorithm 1 NUQLS
Input: number of realizations $S$, weights $\widehat{\bm{\theta}}$.
for $s=1$ to $S$ do
  $\bm{\theta}^{\star}_{s}\leftarrow$ Run (stochastic) GD from $\bm{\theta}_{0,s}$ to
    (approximately) solve ˜3 and obtain $\bm{\theta}^{\star}_{s}$
end for
return $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},.)\}_{s=1}^{S}$
We now present Neural Uncertainty Quantification by Linearized Sampling (NUQLS), our post-hoc sampling method for quantifying the uncertainty of a trained NN. We begin by presenting the motivation and a high-level overview of our method. Subsequently, we provide theoretical justification, demonstrating that, under specific conditions, the NUQLS samples represent draws from the approximate posterior of the neural network, which is equivalent to a GP defined by the NTK.
@@ 3.1 Motivation and High-level Overview
[…]
where ${\mathbf{J}}(.,{\bf x})=[\partial{\bf f}(.,{\bf x})/\partial\bm{\theta}]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\in^{c\times p}$ is the Jacobian of $\bf f$}. Using the linear approximation ˜2, we consider
@@
In overparameterized settings, ˜3 may have infinitely many solutions. To identify these solutions and create our ensemble of linear predictors, we employ (stochastic) gradient descent, initialized at zero-mean isotropic Gaussian perturbations of the trained parameter, $\widehat{\bm{\theta}}$. The pseudo-code for this algorithm is provided in Algorithm˜1. For a given test point $\bf x$$^{\star}$}, the mean prediction and uncertainty can be computed using $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x}^{\star})\}_{s=1}^{S}$.
Note that while the training cost for a linearised network is only slightly higher per epoch compared to standard NN training, each network in the NUQLS ensemble is initialized in a neighborhood of a local minimum of the original NN. As a result, NUQLS often requires significantly fewer epochs to converge, leading to an order-of-magnitude computational speedup relative to DE (see Tables˜1, 2 and 9 for wall-clock time comparisons, and Appendix˜E for a more in-depth analysis of the computation costs.).
@@ 3.2 Theoretical Analysis
We now establish the key property of Algorithm˜1: under mild conditions, NUQLS generates samples from the approximate posterior of the neural network, which in many cases corresponds to a Gaussian process defined by the NTK. Proofs are provided in Appendix˜A.
Suppose $\bm{\theta}^{\ddagger}$ is any solution to ˜3. Using $\bm{\theta}^{\ddagger}$, one can construct a family of solutions to ˜3 as
@@
[…]
@@
  • strictly convex in its first argument, and a solution to ˜3 exists.
The problem ˜5 admits a unique solution.
As it turns out, any solution of the form ˜4 can be efficiently obtained using (stochastic) gradient descent.
@@
Consider the optimization problem ˜3 and assume ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank.
  • (Gradient Descent) Suppose $\ell({\bf f},{\bf y})$ is strictly convex with locally Lipschitz continuous gradient, both with respect to $\bf f$}, and the problem ˜3 admits a solution. Gradient descent, initialized at $\bf z$} and with appropriate learning rate, converges to ˜4.
  • (Stochastic Gradient Descent) Suppose $\ell({\bf f},{\bf y})$ is strongly convex with Lipschitz continuous gradient, both with respect to $\bf f$}, and any solution to the problem ˜3 is interpolating. Stochastic gradient descent, initialized at $\bf z$} and with small enough learning rate, converges to ˜4 with probability one.
We note that the local smoothness requirement in the first part of Theorem˜3.2 is a relatively mild assumption; for example, it holds if we simply assume that $\ell$ is twice continuously differentiable. Also, the full-row rank assumption on ${\mathbf{J}}_{\mathcal{X}}$ in the second part of Theorem˜3.2 is reasonable for highly over-parameterized networks; e.g., see Liu et al. (2022).
Now, suppose ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank and the assumption of Lemma˜3.1 holds, ensuring the existence of the unique solution $\bm{\theta}^{\ddagger}$ to ˜5. Noting ${\textnormal{Range}}\left({\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)={\textnormal{Range}}({\mathbf{J}}_{\mathcal{X}}^{\dagger})$, we can write ˜4 as
@@
where $\bf w$} is a vector for which $\bm{\theta}^{\ddagger}={\mathbf{J}}_{\mathcal{X}}^{\dagger}{\bf w}$, and ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}\triangleq{\mathbf{J}}_{\mathcal{X}}{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}_{\mathcal{X}}={{\mathbf{K}}}_{\widehat{\bm{\theta}}}(\mathcal{X},\mathcal{X})\in^{nc\times nc}$ is the Gram matrix of the empirical NTK ˜1 on the training data $\mathcal{X}$. Setting ${\bf z}=\widehat{\bm{\theta}}-{\bf z}_{0}$ for some ${\bf z}_{0}$, we get